OpenStudy (anonymous):

Explain how you would use the quadratic formula to solve x2 + 5x = -2. Why is the quadratic formula the best method to use?

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + 5x = -2\] \[\Large x^2 + 5x + 2 = 0\] Now use the quadratic formula. It's best to use it because it can solve ANY quadratic equation. \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(5)\pm\sqrt{(5)^2-4(1)(2)}}{2(1)}\] \[\Large x = \frac{-5\pm\sqrt{25-(8)}}{2}\] \[\Large x = \frac{-5\pm\sqrt{17}}{2}\] \[\Large x = \frac{-5+\sqrt{17}}{2} \ \text{or} \ x = \frac{-5-\sqrt{17}}{2}\]